OpenStudy (anonymous):

Solve 3c + 4 = -4c - 17

OpenStudy (diyadiya):

3c+4c=-17-4 7c=-21 c=-3

OpenStudy (anonymous):

3c + 4 = -4c - 17 -4 -4 3c = -4c - 21 +4c +4c 7c = -21 /7 /7 c = -3